The number of ways in which 2160 can be written as a product of factors which are relatively prime

Grade:8

1 Answers

Vikas TU
6 years ago
.Factor of 2160 are 1,2,3,4,5,6,8,9,10,12,15,16,18,20,24,27,30,36,40,45,48,54,60,72,80,90,108,120,135,144,180,216,240,270,360,432,540,720,1080,2160, 
Entirety of elements of 2160 = 7440 
Composing 7440 as far as forces of primes 
7440=2^4*3*5*31 
no. of methods for composing a number as result of two co prime factors=2^(n-1) 
where n=number of prime element of a number 
henceforth 2^(4-1) =8